OpenStudy (anonymous):

Question :12 - Signals: Find whether the given Discrete-time signal is Periodic. If so, then find its Periodicity.. \[\huge \color{green}{ x[n] = e^{\frac{j \cdot \pi \cdot n}{16}} \cos(\frac{n \cdot \pi}{17})}\]

OpenStudy (foolaroundmath):

It will definitely be periodic the exponent is periodic in \(32 \pi \) and the \(\cos\) is periodic in \(34 \pi\). So \(x[n]\) is periodic in \(\pi\cdot LCM(32,34) = 544 \pi\) Whether there exists a smaller period, I don't want to find out.

OpenStudy (anonymous):

How can exponential has period of \(32 \pi\) ?? It is discrete-time signal and not Continuous-time.. :)

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@iambatman save me.. :)

OpenStudy (anonymous):

The Complex Exponential is Periodic with N = 32 Samples.. And the cosine term is Periodic with N = 34 Samples.. But what will be Overall Period of this overall sequence?? LCM of (32, 34) ??

OpenStudy (unklerhaukus):

for a simple signal \[x(t)=A\cos(\omega t-\varphi)\] the period is \[T=2\pi/\omega\]

OpenStudy (anonymous):

Uncle Rocks, yes you are right, but this is Discrete Time Signal having sampled values of continuous time signal, n here is an integer..

OpenStudy (anonymous):

\(x[n] = Acos(\Omega n - \phi)\) the period of this sequence is: \[N = \frac{2 \pi}{\Omega} \times r\]

OpenStudy (anonymous):

Where \(r\) is the smallest integer for which N turns to be an Integer.. :)

OpenStudy (anonymous):

Like: \(x[n] = cos(3 \pi n)\) \[N = \frac{2 \pi}{3 \pi } \times r \implies N = \frac{2}{3} \times r\] But here N is not integer, so for smallest, r = 3, N = 2 which becomes Integer..

OpenStudy (anonymous):

So, for this Discrete Signal, N = 2 Samples is the period..

OpenStudy (anonymous):

If you can use any software, then can you show the waveform of my original question?

OpenStudy (unklerhaukus):

measuring on my graph period is = 16.5 which isn't an integer, but it is between 16 and 17

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the complex exponent, recollections of trig conversions to e ... can we convert this into all trig or all exponent?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

By Euler Identity,we can do that..

OpenStudy (amistre64):

it might be easier to see when its all in one form ...

OpenStudy (anonymous):

\[x[n] = [\cos(\frac{ \pi n}{16}) + j \sin(\frac{\pi n}{16})] \cos(\frac{\pi n}{17})\]

OpenStudy (anonymous):

Should I distribute that cosine term into brackets??

OpenStudy (amistre64):

and i recall stuff like: cos(a+b) = cos(a)cos(b) - sin(a)sin(b) cos(a-b) = cos(a)cos(b) + sin(a)sin(b) adding we get cos(a+b) + cos(a-b) = 2cos(a)cos(b) cos(a)cos(b) = cos(a+b)/2 + cos(a-b)/2 not sure how the i parts pan out tho

OpenStudy (amistre64):

i would, to attempt to get this into some linear format

OpenStudy (amistre64):

not sure if theres a thrm or something since i havent taken any courses with this stuff

OpenStudy (anonymous):

I can distribute and then with trig identities I can make them into addition form too.. But will that be helpful? I don't know when it comes to imaginary.. If it were single imaginary only, then also I can find it,, but extra cos term is bothering me..

OpenStudy (amistre64):

sin(a+b) = sin(a)cos(b) + sin(b)cos(a) sin(a-b) = sin(a)cos(b) - sin(b)cos(a) works the same, the j part is just the j part .... i believe the combined trig argument is what i would be looking at

OpenStudy (anonymous):

Then let me simplify to look more clear form of it..

OpenStudy (anonymous):

But it will be painful to do that.. :(

OpenStudy (amistre64):

math is often painful, and hardly ever worth it ... but death absolves all absolutes :)

OpenStudy (anonymous):

Real part goes like this : \[\frac{1}{2}\cos( \pi n \frac{33}{272}) + \frac{1}{2}\cos(\pi n \frac{1}{272})\]

OpenStudy (anonymous):

Right??

OpenStudy (amistre64):

mcd wifi is slow ...

OpenStudy (anonymous):

Imaginary part will be: \[\frac{1}{2}\sin( \pi n \frac{33}{272}) + \frac{1}{2}\sin(\pi n \frac{1}{272})\]

OpenStudy (anonymous):

mcd??

OpenStudy (amistre64):

mcdonalds

OpenStudy (anonymous):

Are you there??

OpenStudy (anonymous):

Or it is your wifi connection from there??

OpenStudy (amistre64):

daughter is working here, and so today i dropped her off and am using the wifi

OpenStudy (amistre64):

but yes, those are fine 33/272 and 1/272

OpenStudy (anonymous):

That's great.. :)

OpenStudy (amistre64):

so what are the 2 periods for those functions? and can you think of a way to combine them so that they work in 1 period togehter?

OpenStudy (anonymous):

I try..

OpenStudy (anonymous):

How about finding \(\omega\) by using : \[\omega = \frac{HCF(Numerator)}{LCM(Denominator)}\]

OpenStudy (amistre64):

if that works, give it a shot :)

OpenStudy (anonymous):

If I find individual periods, then for Discrete time, their sum will be periodic, it is a proven thing..

OpenStudy (anonymous):

For Individual Signals, I got N = 544 Samples..

OpenStudy (anonymous):

\[N = \frac{2 \pi}{\frac{\pi}{272}} \implies N = 544\]

OpenStudy (amistre64):

thats what im getting

OpenStudy (anonymous):

This is Individual periods..

OpenStudy (anonymous):

The common \(\omega\) will come out be : \[\omega = \frac{HCF(\pi , 33 \pi)}{LCM(272, 272)} \implies \omega = \frac{\pi }{272}\]

OpenStudy (anonymous):

This will give N = 544 Samples..

OpenStudy (anonymous):

I was just about to say to you that please check on wolfram.. :)

OpenStudy (amistre64):

theres prolly a simpler way to go about it, but i dont do simple lol

OpenStudy (anonymous):

yes, @amistre64 thank you so much.. :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

What was the simpler method, I don't get yet.. :)

OpenStudy (anonymous):

You have more time??

OpenStudy (anonymous):

Even this looks a simpler method now, atleast it is giving me an answer. :P

OpenStudy (amistre64):

simpler method, find someone smarter than me ;) someones asking me something in another question, so ill be around

OpenStudy (anonymous):

I have one doubt regarding limits of integration, can you clarify?

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} x(\tau) h(t - \tau) d \tau\]

OpenStudy (anonymous):

Here, if I put \(t - \tau = \lambda\) Then how the limits will change?

OpenStudy (anonymous):

Still -infinity to infinity or infinity to -infinity?? The initials are the limits of \(\tau\)..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Tell me about the limits..

OpenStudy (anonymous):

Reading for so long..

ganeshie8 (ganeshie8):

\[\int\limits_{-\infty}^{\infty} x(\tau) h(t - \tau) d \tau = \int\limits_{\infty}^{-\infty} x(t-\lambda ) h(\lambda ) (-d\lambda) = \int\limits_{-\infty}^{\infty} x(t-\lambda ) h(\lambda ) d\lambda \]

ganeshie8 (ganeshie8):

have to cuz you always ask hard questions :o

OpenStudy (anonymous):

Is this hard?

OpenStudy (anonymous):

Oh my God, what the hell, why I did not see - with d(lambda) too.. :) I am just a useless guy..

OpenStudy (anonymous):

Wait, how this goes if I have summation in place of integration?

OpenStudy (anonymous):

\[\sum_{k = -\infty}^{\infty} x[k] h[n-k]\]

OpenStudy (anonymous):

If here I put : \(n - k = p\) ??

ganeshie8 (ganeshie8):

how do we define \(\large \sum \limits_{k=0}^{-10}f(k)\) ?

OpenStudy (anonymous):

Oh!! I think in discrete summation, limits can be interchanged without any worry..

OpenStudy (anonymous):

As it is a Sum, it can be written from left to right or right to left..

OpenStudy (anonymous):

There -10 will come beneath and 0 will go upwards..

OpenStudy (anonymous):

\[\sum_{k = -10}^{0} f(k)\]

ganeshie8 (ganeshie8):

it is better to write it in below form : \[\large \sum \limits_{-10\le k\le 0}f(k)\]

OpenStudy (anonymous):

So: \[\sum_{k = \infty}^{-\infty} = \sum_{k =- \infty}^{\infty} \]

OpenStudy (anonymous):

Okay, then??

ganeshie8 (ganeshie8):

\[\large \sum_{ -\infty \lt k \lt \infty } x[k] h[n-k] \]

ganeshie8 (ganeshie8):

changing variables should not have any effect on limits i guess

OpenStudy (anonymous):

We can't write it like this?? \[\large \sum_{ -\infty \lt k \lt \infty } x[k] h[n-k] = \sum_{k = -\infty}^{\infty} x[k]h[n-k]\]

OpenStudy (anonymous):

Yes, limits will not change in case of summation..

ganeshie8 (ganeshie8):

i don't see anything wrong, but i have never seen a negative infinity as lower limit in a sum

OpenStudy (anonymous):

You have not seen??

OpenStudy (anonymous):

I am seeing it the time I begin Convolution of Discrete Signals.. :)

ganeshie8 (ganeshie8):

it means you're adding up area below the funtion from left end to right end

OpenStudy (anonymous):

I am always thankful to you @ganeshie8 , thanks for your help here.. For about 20 minutes, I was thinking, how I can change the limits again to -infty to infty..

OpenStudy (anonymous):

Thank you.. :)